# Present address: ]TRIUMF, Vancouver, BC V6T 2A3, Canada

### v2

@@ V Conclusions
In conclusion, we have conducted new, high-precision mass measurements of 129g,mSn, 131g,mSn, and 132g,mSb using the Canadian Penning Trap at Argonne National Laboratory’s ATLAS user facility’s CARIBU facility. Our results for these ground and isomeric states are in good agreement with most of the preceding literature values, and provides the highest-precision determination of the excitation energies of these long-lived isomeric states. We then used these new excitation energies to determine the effective thermal transitions between the ground and isomeric states, their relative $\beta$ decay rates, and the associated cross sections into and out of these states. The excitation energies were also used to estimate the thermalization temperature with respect to $\beta$-decay, above which these nuclei may be correctly treated in reaction networks as one species. Reaction network calculations were used to establish that 129mSn must be treated as an astromer in the $i$-process and $r$-process. Both 131mSn and 132mSb were determined not to behave as astromers and these isotopes can be safely treated as one species in reaction network calculations. As the thermalization of these nuclei is dependent on the excited states above the isomers, a full understanding of their astromeric nature would benefit from detailed spectroscopy of these states and on the determination of their feeding via $\beta$ decays.

### v3

@@ V Conclusions
In conclusion, we have conducted new, high-precision mass measurements of 129g,mSn, 131g,mSn, and 132g,mSb using the Canadian Penning Trap at Argonne National Laboratory’s ATLAS user facility’s CARIBU facility. Our results for these ground and isomeric states are in good agreement with most of the preceding literature values, and provide the highest-precision determination of the excitation energies of these long-lived isomeric states directly from their masses. We then used these new excitation energies to determine the effective thermal transitions between the ground and isomeric states, their relative $\beta$ decay rates, and the associated cross sections into and out of these states. The excitation energies were also used to estimate the thermalization temperature with respect to $\beta$-decay, above which these nuclei may be correctly treated in reaction networks as one species. Reaction network calculations were used to establish that 129mSn must be treated as an astromer in the $i$-process and $r$-process. The excited state in 131Sn may need to be treated as an astromer depending on the astrophysical conditions. The state in 132Sb was determined not to behave as astromers under the conditions studied in this work. As the thermalization of these nuclei is dependent on the excited states above the isomers, a full understanding of their astromeric nature would benefit from detailed spectroscopic studies of the spins and parities of the known states, a search for potential missing states, and on the determination of their feeding via $\beta$ decays.